Maria is eight years old than precy ten years ago, the product of their age is 84. How to find the present?
Answer by ankor@dixie-net.com(22740)

let m = M's age
let p = P's age
:
write an equation for each statement
:
Maria is eight years older than precy
m = p + 8
ten years ago, the product of their ages was 84.
(m-10)*(p-10) = 84
Replace m with (p+8)
((p+8)-10)*(p-10) = 84
(p-2)*(p-10) = 84
FOIL
p^2 - 10p - 2p + 20 - 84 = 0
a quadratic equation
p^2 - 12p - 64 = 0
Factors to
(p+4)(p-16) = 0
two solutions, but we only want the positive solution
p = 16 yr is P's age
then
16 + 8 = 24 yrs is M's age
:
:
See if that checks out in the statement
"10 years ago, the product of their ages was 84."
(16-10)*(24-10) = 84